OpenStudy (anonymous):

√¯32 + 2/√¯2 - √¯2/9 How to simplify?

OpenStudy (campbell_st):

well you need to recognise 32 = 16 x 2 so then \[\sqrt{32} = \sqrt{16 \times 2} = \sqrt{16} \times \sqrt{2}\] so you need to simplify this... then you have 3 like terms that can be added and subtracted.... hope it helps

OpenStudy (anonymous):

Yeah i know this. 4√¯2 But what about others?

OpenStudy (trojanpoem):

As champbell_st said: \[\sqrt{\frac{ 2 }{ 9 }} = \frac{ \sqrt{2} }{ 3 } = \frac{ 1 }{ 3 }\sqrt{2}\] 4sqrt(2) + 2sqrt(2) - 1/3 sqrt(2) sqrt(2) [ 4 + 2 - 1/3] sqrt(2) [ 12/3 + 6/3 - 1/3] sqrt(2) [ 12 + 6 - 1/3] sqrt(2) [17/3]

OpenStudy (anonymous):

You made one mistake.. But i did it right so thank u for your directions! http://i68.tinypic.com/140b61e.jpg